The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
4/25

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

4 / 25
Totally there are 3 groups of equivalent fractions.
1 / 6 = (a), (e), (h), (j), (k)
1 / 5 = (b), (f), (g)
4 / 25 = (d), (i), (l)

​Thanks